The following data relate to Webster Manufacturing:

Anticipated sales...............................     50,000 units

Normal sales price/unit.....................     $20

Variable manufacturing cost/unit..... $6

Variable selling cost/unit.................... $2

Fixed manufacturing cost.................     $150,000 ($3.00/unit)

Fixed selling cost ............................ $200,000 ($4.00/unit)

Total cost/unit (full absorption)........     $15.00

Webster has the opportunity to take a one-time special order from a customer who wants to buy 10,000 units, but is only willing to pay $14.00 per unit. The special order will not affect Webster’s anticipated sales to other customers or the company’s fixed costs.

Should Webster accept the special order? Why or why not? Show your calculations.
Answer #1
As the special order will not affect the existing
market, the decision will be based on incremental
revenues and costs associated with the special order.
Incremental revenue = 10000*14 = $   1,40,000
Incremental costs = 10000*(6+2) = $       80,000
Incremental operating profit $       60,000